(x-1)^5+x^5=0 value of x
? ? ? ? ? ? ? ? ? ? ? ? ? ? ? ? ? ? ? ? ? ? ? ?  ? ? ~ ? ? ? ? ? ? ? ? ? ~ ? ? ? ? ~ ~ ~ ? 

Grade:12

3 Answers

Aditya Gupta
4 years ago
hello rahul, write it as
(x-1)^5= – x^5= ( – x)^5
so, taking fifth roots both sides,
x – 1= – x
or x= ½ = 0.5
kindly approve :)
Arun
4 years ago
Dear student
 
As we see that x -1 = - x
Hence
 
2x = 1
x = ½
 
Hope it helps
 
Regards
Arun
 
In case of any difficulty, please feel free to ask.
MANTHAN RAO
4 years ago
ANSWER)1/2
EXPLAINATION)
(x-1)^5+x^5=0
+X^5AN BE MOVE ON RHS AS PER EQUATION RULES
ALSO THE SIGN OF +X^5WIL ALSO CHANGE THAT IS IT WILL BECOME -X^5
(X-1)^5=-X^5
WE CAN TAKE FIFTH ROOT N BOOTH SIDE
THEREFORE
X-1=-X
X+X=1
AS X AND X ARE SAME VARIABLE HENCE THEY CAN BE ADDED
2X=1
X=1/2 ANS.